2018 2017 Assets Cash and equivalents $14,000 11,000 Accounts receivable 35,000 30,000 Inventories 30,120 27,000 Total current assets 79,120 68,000 Net plant and equipment 48,000 45,000 Total assets $127,120 $113,000 Liabilities and Equity Accounts payable 8,500 10,500 Accruals 7,600 6,000 Notes payable 6,100 5,150 $ 19,650 Total current liabilities 24,200 Long-term bonds 10,000 10,000 $29,650 Total liabilities 34,200 Common stock (4,000 shares) 50,000 50,000 Retained earnings 42,920 33,350 Common equity 92,920 83,350 Total liabilities and equity $127,120 $113,000 Income Statement for Year Ending December 31, 2018 Sales $245,000 Operating costs excluding depreciation and amortization 200,000 ЕBITDA $45,000 Depreciation & amortization 5,000 EBIT $40,000 Interest 3,750 EBT 36,250 Taxes (40%) 14,500 Net income $21,750 Dividends paid 12,180
Enter your answers in millions. For example, an answer of $25,000,000,000 should be entered as 25,000. Round your answers to the nearest whole number, if necessary. a. What was net operating working capital for 2017 and 2018? Assume that all cash is excess cash; i.e., this cash is not needed for operating purposes. 2017 $ million 2018 $ million b. What was Arlington's 2018 free cash flow? million c. Construct Arlington's 2018 statement of stockholders' equity. Statement of Stockholders' Equity, 2018 Total Stockholders Retained Common Stock Earnings Equity Amount Shares Balances, 12/31/17 million million million million 2018 Net Income million Cash Dividends million Addition to retained earnings million million Balances, 12/31/18 million million million d. What was Arlington's 2018 EVA? Assume that its after-tax cost of capital is 10 % . Round your answer to two decimal places. million e. What was Arlington's MVA at year-end 2018? Assume that its stock price at December 31, 2018 was $25. Round your answer to two decimal places million
Answer #1

A. NOWC for 2017 is $53500

NOWC for 2018 is $61020

NOWC= (cash+inventory+AR) - (accruals + AP)

B. FCF= EBIT-tax + dep&A -CAPEX - change in WC

= $11830

C. for year 2017- SHARES 4000 AMOUNT $50,000 RETAINED EARNINGS $33350 shareholder's equity (shares+retained earnings) =$83350

net income for year 2018 is $21750

cash dividend paid is $12180

addition to retained earnings (42920-33350) =$9570

balance for year 2018 shares 4000 amount $50,000 retained earnings $42920 shareholder's equity $92920

D. EVA = NET INCOME - COST OF EQUITY CAPITAL

= 21750-(10% * 50000) = $16750

E. MVA= MARKET VALUE OF STOCK- TOTAL COMMON EQUITY

= (25*4000) - 92920 = $7080